Cash Company has income from operations of $31,584, invested assets of $141,000, and sales of $394,800. Use the DuPont formula to compute the return on investment. If required, round your answers to two decimal places.

a. Profit margin %
b. Investment turnover
c. Return on investment %
Answer #1

Margin = Net operating income/Sales

= 31,584/394,800

= 8%

Turnover = Sales/Average operating assets

= 394,800/141,000

= 2.80

Return on investment = 8*2.80

= 22.40%